THE PUBLICLY AVAILABLE PREBIOTIC INTERSTELLAR MOLECULAR SURVEY (PRIMOS): EXPANDING SPECTROSCOPIC CHARACTERIZATIONS, EXTENDING TO NEW SOURCES, AND ADDING TO THE KNOWN MOLECULAR INVENTORY

\underline{BRETT A. McGUIRE}, P. BRANDON CARROLL, Division of Chemistry and Chemical Engineering, California Institute of Technology, Pasadena, CA 91125; JOANNA F. CORBY, Department of Astronomy, University of Virginia, Charlottesville, VA 22904; RYAN A. LOOMIS, Department of Chemistry, University of Virginia, Charlottesville, VA 22904; GEOFFREY A. BLAKE, Division of Chemistry and Chemical Engineering and Division of Geological and Planetary Sciences, California Institute of Technology, Pasadena, CA 91125; JAN M. HOLLIS, NASA Goddard Space Flight Center, Greenbelt, MD 20771; FRANK J. LOVAS, National Institute of Standards and Technology, Gaithersburg, MD 20899; PHILIP R. JEWELL, and ANTHONY J. REMIJAN, National Radio Astronomy Observatory, Charlottesville, VA 22903.

The publicly available Green Bank Telescope PRebiotic Interstellar MOlecular Survey (PRIMOS) conducted towards Sgr B2(N) provides high resolution, high-sensitivity observations with near-continuous frequency coverage from ∼300 MHz - 50 GHz. Of the eleven new molecular detections in the last year, five (45%) are a direct result of the PRIMOS observations. Further, these observations have recently been used to detect the predicted, but previously unobserved, J = 1−0 and J = 2−1 transitions of the newly detected l-C_{3}H^{+} ion. Here, we discuss the analysis of these transitions, as well as recent work to extend the PRIMOS observations to three new regions of interest: VY Canis Majoris, IRC+10216, and NGC 2023. Finally, we highlight the utility of cm-wave surveys in new molecular detections, as well as the value of publicly-available surveys in the approaching era of data-enabled, analysis-limited astrochemistry.